The following table shows the annualized return of S&P total return index from 1999 to 2003:

Year Total Return (%)

1999 21.0

2000 -9.1

2001 -11.9

2002 -22.1

2003 28.7

(a) Calculate the geometric return from 1999 to 2003.

(b) Calculate the mean absolute deviation and the sample standard deviation of the index

return.

(c) Calculate the 45th-percentile of the index return.

(d) Sketch the drawdown curve. What is the maximum drawdown of the index return over this

period?

(e) Assume an annualized risk-free rate of 5%, what is the Sharpe ratio of the index return over

this period?
Answer #1

(a) Geometric return is calculated by solving the following equation:

Returneo = V(1+ri) x (1+r2) x (1 + r3) x (1 + r) x (1 + r) - 1

Return geo = V(1+0.21) (1 - 0.091) x (1 -0.119) (1 - 0.221) X (1+0.287) -

Returngeo = -0.57669%

(b) Mean absolute deviation is calculated below:

Year Returns MAD
1999 21.0% abs(0.21-0.0132)= 19.68%
2000 -9.1% abs(-0.091-0.0132)= 10.42%
2001 -11.9% abs(-0.119-0.0132)= 13.22%
2002 -22.1% abs(-0.221-0.0132)= 23.42%
2003 28.7% abs(0.287-0.0132)= 27.38%
Sum 6.6% Sum 94.1%
Count 5 Count 5
Mean 6.6/5= 1.32% MAD 94.1/5 = 18.82%

Sample standard deviation is calculated below:

Year Returns Squared deviations
1999 21.0% (0.21-0.0132)^2= 3.87%
2000 -9.1% (-0.091-0.0132)^2= 1.09%
2001 -11.9% (-0.119-0.0132)^2= 1.75%
2002 -22.1% (-0.221-0.0132)^2= 5.48%
2003 28.7% (0.287-0.0132)^2= 7.50%
Sum 6.6% Sum 19.7%
Count 5 Count 4
Mean 0.0132 Variance 19.7/4 = 4.92%
Sample standard deviation 4.92^0.5 = 22.19%
  • As we need to find the sample standard deviation, we need to divide the sum of the squared deviations by n-1 = 5-1 = 4 to find the variance
  • After calculating the variance, we need to take a square-root of it to find the sample standard deviation

(c)

  • To calculate the 45th percentile we first need to reorder the returns in ascending order as follows:
Year Returns
2002 -22.1%
2001 -11.9%
2000 -9.1%
1999 21.0%
2003 28.7%
  • Next we multiply the required percentile with the number of observations: 0.45 x 5 = 2.25
  • Next we round off 2.25 to the next highest whole number ie 3
  • The third observation is the 45th percentile , which is 9.1%

(d) Drawdown is calculated below:

Year Returns Drawdown
1999 21.0% 0.731707317
2000 -9.1% -0.317073171
2001 -11.9% -0.414634146
2002 -22.1% -0.770034843
2003 28.7% 1
  • Maximum return is 28.7%
  • We divide each return with this to find the draw down of the year or what percentage is it of the maximum return
  • Then we plot the same as follows:
  • Drawdown 20b0 2001 20b2 2003 -0.21999 -0.4 -0.5

We can see the lowest fall is in 2002 when the return goes down to -22.1% and the drawdown is -0.770034843 ie the maximum drawdown

(e)

Sharpe ratio = Retur ngeo - Retur riskfree standard deviation

-0.57669% - 5% Sharpe ratio = - -=-1.133007667 4.92%